OpenStudy (iwanttogotostanford):

HELP PLZ: Simplify 2 times square root 12 minus 2 square root 9 plus square root of 112.

OpenStudy (iwanttogotostanford):

I got the last one btw

OpenStudy (studygurl14):

\(\Large 4\sqrt{24}-4\sqrt{8}+\sqrt{98}\)

OpenStudy (iwanttogotostanford):

yes @StudyGurl14

OpenStudy (studygurl14):

okay, so do you know how to simplify radicals?

OpenStudy (iwanttogotostanford):

kind of but for this one i forgot.

OpenStudy (studygurl14):

okay, let's start with 8. Try seperating it into its factors

OpenStudy (iwanttogotostanford):

4 and 2?

OpenStudy (studygurl14):

yep. keep going further. break down the 4

OpenStudy (iwanttogotostanford):

so break the 4 to 2 and 2

OpenStudy (studygurl14):

yep, so 8 = 2*2*2 Now, watch the magic: \(\Large\sqrt{8}=\sqrt{2\times2\times2}=\sqrt{2^2\times2}=2\sqrt{2}\)

OpenStudy (studygurl14):

do you see how that worked?

OpenStudy (iwanttogotostanford):

yeah !

OpenStudy (iwanttogotostanford):

that actually did help a lot thanks!

OpenStudy (studygurl14):

actually, I just realized when didn't even have to break down the 4, because I'm sure you already know \(\sqrt{4}=2\)

OpenStudy (iwanttogotostanford):

yes

OpenStudy (studygurl14):

Yep, and since we had \(\Large-4\sqrt{8}\)...\(\Large-4(2\sqrt{2}=-8\sqrt{2}\) Now, do you think you could do the same with \(\Large 4\sqrt{24}\)?

OpenStudy (iwanttogotostanford):

yes, so do I simplify inside the radical first? or the 4.

OpenStudy (iwanttogotostanford):

@StudyGurl14

OpenStudy (studygurl14):

inside the radical

OpenStudy (iwanttogotostanford):

@StudyGurl14 so i would get 6 and 4

OpenStudy (iwanttogotostanford):

@StudyGurl14

OpenStudy (studygurl14):

yep, so you'd get \(\Large 4\sqrt{24}=4\sqrt{6\times4}=?\)

OpenStudy (iwanttogotostanford):

would I factor out the 6 or the 4

OpenStudy (studygurl14):

the 4, because it can easily be turned into 2

OpenStudy (iwanttogotostanford):

oh ok! so I would now have 8 on the outside of the radical correct?

OpenStudy (studygurl14):

yep! So, so far we have:\(\Large 8\sqrt{6}-8sqrt{2}+\sqrt{98}\) Can you do the last one now?

OpenStudy (iwanttogotostanford):

how did u get the 98

OpenStudy (studygurl14):

that's part of the original equation, remember? We haven't simplified that one yet

OpenStudy (iwanttogotostanford):

oh yea

OpenStudy (iwanttogotostanford):

i can't think off the top of my head what goes into 98....

OpenStudy (studygurl14):

well...it's an even number right? so try dividing it by 2

OpenStudy (studygurl14):

tell me what you get

OpenStudy (iwanttogotostanford):

49!

OpenStudy (studygurl14):

right! and what's \(\sqrt{49}\)?

OpenStudy (iwanttogotostanford):

8 would be the factor!

OpenStudy (iwanttogotostanford):

wait no, 7 sorry

OpenStudy (studygurl14):

:) right. so can you finish simplifying that one?

OpenStudy (iwanttogotostanford):

well, i would have 7 on the outside of the radical and then 7 on the inside right? but then what?

OpenStudy (studygurl14):

no, almost \(\Large\sqrt{98}=\sqrt{49\times 2}=7\sqrt{2}\)

OpenStudy (iwanttogotostanford):

im a little confused on this last step, i really want to learn but just still a lil confused

OpenStudy (studygurl14):

What exactly are you confused about?

OpenStudy (iwanttogotostanford):

well, what to do exactly next or how to know what to do next

OpenStudy (studygurl14):

Oh, okay. So now we have \(\Large 8\sqrt{6}-8\sqrt{2}+7\sqrt{2}\) right?

OpenStudy (iwanttogotostanford):

yes

OpenStudy (studygurl14):

Okay, so notice how both the -8 and 7 have a sqrt 2?

OpenStudy (iwanttogotostanford):

yep!

OpenStudy (studygurl14):

You can combine them like normal numbers. What is -8 + 7?

OpenStudy (iwanttogotostanford):

-1!

OpenStudy (studygurl14):

yep, so you get \(\Large 8\sqrt{6}-\sqrt{2}\) You're done

OpenStudy (iwanttogotostanford):

oh wow, ok so what happened to the 98 again?

OpenStudy (studygurl14):

Remember, 98 = 49 times 2, so...\(\Large\sqrt{98}=\sqrt{49\times 2}=\sqrt{49}\times\sqrt{2}=7\sqrt{2}\)

OpenStudy (iwanttogotostanford):

and could i please have your help on one more?? you were super helpful and i actually learned

OpenStudy (studygurl14):

Make sense?

OpenStudy (iwanttogotostanford):

yes, completely!

OpenStudy (studygurl14):

Great. :) and yes, I'll help you with another, but in a new thread, since this one is awfully long now

OpenStudy (iwanttogotostanford):

ok, thank you!